Problem 8-05A a-b (Video) The bank portion of the bank reconciliation for Bonita Company at November 30, 2020, was as followsThe cash records per books for December showed the following. Date 12-1 12-2 12-2 12-4 12-8 Number 3475 3476 3477 3478 3479 3Using the four steps in the reconciliation procedure, prepare a bank reconciliation at December 31. (Round answers to 2 decimSHOW LIST OF ACCOUNTS LINK TO TEXT Prepare the adjusting entries based on the reconciliation. (Credit account titles are autoPlease try to keep the same format!
Answer #1
BONITA COMPANY
Bank Reconciliation
December 31, 2022
Cash balance per bank statement 19645.80
Add: Deposits in transit 1690.40
21336.20
Less: Outstanding checks
No. 3470 720.10
No. 3474 1051.30
No. 3478 621.30
No. 3481 807.40
No. 3484 796.70
No. 3486 889.50
4886.30
Adjusted cash balance per bank 16449.90
Cash balance per books 14395.20
Add: Electronic funds transfer 2420.00
16815.20
Less:
Error in recording check 270.00
Error in deposit 9.00
Check printing charge 86.30
365.30
Adjusted cash balance per books 16449.90
2
Dec.31     Cash 2420.00
                         Accounts Receivable 2420.00
(To record electronic funds transfer received by bank)
Dec 31     Bank service charge expense/Miscellaneous expense 86.30
                         Cash 86.30
(To record bank service charge)
Dec.31     Accounts Payable 270.00
                         Cash 270.00
(To correct error in recording check)
Dec.31     Accounts Receivable 9.00
                         Cash 9.00
(To correct error in deposit)
Note:
Error in recording check is for check 3485. Error amount = 520.80-250.80 = $270
Error in deposit is for 12/20 deposit. Error amount = 2954-2945 = $9